Mrs. Robinson plays bingo on two nights and wins a total of 25 games in both nights. She however won 7 more games on the first night than the second night. How many games did she win on each night?
Found 2 solutions by checkley77, amoresroy:
Answer by checkley77(12569)

X+X+7=25
2X=25-7
2X=18
X=18/2
X=9 GAMES WON THE SECOND NIGHT
9+7=16 GAMES WON THE FIRST NIGHT.
PROOF:
9+16=25
25=25

Answer by amoresroy(333)

Let x = no. of games won on 1st nignt,
x + x - 7 = 25
2x = 25 + 7
x = 32
x = 16
2nd night
16-7 = 9